What are the unit of the rate constant for a 1/2-order reaction? express all concentrations in molarity and time in seconds.
1
Expert's answer
2020-09-14T08:19:45-0400

Rate = -dCA/dt = KCAn

mole L-1 sec-1 = K(mole L-1) n

n= order of reaction

unit of K= mol1-n Ln-1 sec-1

For ½ order reaction n=1/2

Unit of K = mol1/2 L-1/2 sec-1 [molarity(M)= mol/L]

               =M1/2 sec-1(Answer)